Solve by the substitution method
8m + n = 5
m - 9n = 28

Answer by mananth(16949)

8m + n = 5
m - 9n = 28
n = 5-8m
plug the value of n in the equation below
m - 9n = 28
m-9(5-8m)=28
m-45+72m=28
73m-45=28
add 45 to both sides-m-45+45=28+45
73m=73
m=1
plug the value of m in the equation n=5-8m
n=5-8
n=-3